# Shadow and Polarization Images of Rotating Black Holes in Kalb-Ramond Gravity Illuminated by Several Thick Accretion Disks

### v1

@@ 6.4 Polarization Images Illuminated by HOU Disk
Figure 16: Stokes parameters $\tilde{I}_{o},\tilde{Q}_{o},\tilde{U}_{o},\tilde{V}_{o}$ in the HOU disk model. The fixed parameters are $a=0.1,\varpi=0.3,\varsigma=0.69,\theta=65^{\circ}$.
@@
Figure 17: Polarized images of black hole shadow in the HOU disk model. The accretion flow follows the infalling motion, with fixed parameters $a=0.1,\theta=65^{\circ}$.
@@ 7 Conclusion and Discussion
This paper investigates the optical and polarization imaging characteristics of rotating black holes in KR Gravity under thick accretion disk models. First, we reviewed the basic properties of KR black holes and provided definitions for null geodesics and the photon sphere. Then, we considered two thick accretion disk models: the phenomenological model and the HOU disk model. By numerically solving the geodesic equations and the radiative transfer equations, we obtained the corresponding black hole optical images and polarization structures.
@@
For the phenomenological model, we considered both isotropic and anisotropic radiation scenarios. In the case of isotropic radiation, we fixed the accretion flow motion to the infalling motion, with an observation frequency of $230\,\mathrm{GHz}$. The results show that increasing the spin parameter $a$ enhances the coordinate dragging effect, making the intensity on the left side of the higher-order images significantly greater than on the right. The increase in the spontaneous Lorentz violating parameters $\varsigma$ and $\varpi$ reduces the overall size of the higher-order images, but has little effect on their shape. Notably, unlike the geometrically thin accretion disk, at high observation inclination angles, the inner shadow in the thick disk model is obscured by radiation from outside the equatorial plane, causing the central dark region to split into two parts. The impact of $(a,\varsigma,\varpi)$ on the anisotropic radiation images follows a similar trend to that of isotropic radiation, but in the anisotropic case, the higher-order images are stretched vertically, forming an elliptical shape that reflects the geometric properties of the magnetic field structure. To match the limited resolution of EHT and to more intuitively display the intensity distribution, we further plotted the blurred images and intensity distributions along the xx–axis and yy–axis. The results show that under Gaussian blurring, the distinguishability of the event horizon’s outline is reduced, while the intensity distribution clearly reflects the varying of the size of higher-order image with model dependent parameters.
@@
For the HOU disk model, we found that the distribution of higher-order images is narrower compared to the phenomenological model, and the effect of radiation from outside of the equatorial plane on the inner shadow is significantly reduced. This is because, for specific parameters, the HOU disk is geometrically thinner in some regions under the conical approximation. For polarization images, the linear polarization is negatively correlated with the intensity, and the minimum polarization degree occurs near the higher-order images. The magnitude and direction of the polarization vector are highly sensitive to $(\varsigma,\varpi)$, indicating that the polarization characteristics of rotating KR black holes can effectively reflect their intrinsic spacetime structure. Interestingly, unlike the thin accretion disk model, in the thick disk model, the gravitational lensing effect of radiation from outside of the equatorial plane causes polarization vectors to spread across the entire image plane, leading to more complex polarization patterns.

### v2

@@ 6.4 Polarization Images Illuminated by BAAF Disk
Figure 16: Stokes parameters $\tilde{I}_{o},\tilde{Q}_{o},\tilde{U}_{o},\tilde{V}_{o}$ in the BAAF model. The fixed parameters are $a=0.1,\varpi=0.3,\varsigma=0.69,\theta=65^{\circ}$.
@@
Figure 17: Polarized images of black hole shadow in the BAAF model. The accretion flow follows the infalling motion, with fixed parameters $a=0.1,\theta=65^{\circ}$.
@@ 7 Conclusion and Discussion
This paper investigates the optical and polarization imaging characteristics of rotating black holes in KR Gravity under thick accretion disk models. First, we reviewed the basic properties of KR black holes and provided definitions for null geodesics and the photon sphere. Then, we considered two thick accretion disk models: the RIAF model and the BAAF model. By numerically solving the geodesic equations and the radiative transfer equations, we obtained the corresponding black hole optical images and polarization structures.
@@
For the RIAF model, we considered both isotropic and anisotropic radiation scenarios. In the case of isotropic radiation, we fixed the accretion flow motion to the infalling motion, with an observation frequency of $230\,\mathrm{GHz}$. The results show that increasing the spin parameter $a$ enhances the coordinate dragging effect, making the intensity on the left side of the higher-order images significantly greater than on the right. The increase in the spontaneous Lorentz violating parameters $\varsigma$ and $\varpi$ reduces the overall size of the higher-order images, but has little effect on their shape. Notably, unlike the geometrically thin accretion disk, at high observation inclination angles, the inner shadow in the thick disk model is obscured by radiation from outside the equatorial plane, causing the central dark region to split into two parts. The impact of $(a,\varsigma,\varpi)$ on the anisotropic radiation images follows a similar trend to that of isotropic radiation, but in the anisotropic case, the higher-order images are stretched vertically, forming an elliptical shape that reflects the geometric properties of the magnetic field structure. To match the limited resolution of EHT and to more intuitively display the intensity distribution, we further plotted the blurred images and intensity distributions along the xx–axis and yy–axis. The results show that under Gaussian blurring, the distinguishability of the event horizon’s outline is reduced, while the intensity distribution clearly reflects the varying of the size of higher-order image with model dependent parameters.
@@
For the BAAF model, we found that the distribution of higher-order images is narrower compared to the RIAF model, and the effect of radiation from outside of the equatorial plane on the inner shadow is significantly reduced. This is because, for specific parameters, the BAAF is geometrically thinner in some regions under the conical approximation. For polarization images, the linear polarization is negatively correlated with the intensity, and the minimum polarization degree occurs near the higher-order images. The magnitude and direction of the polarization vector are highly sensitive to $(\varsigma,\varpi)$, indicating that the polarization characteristics of rotating KR black holes can effectively reflect their intrinsic spacetime structure. Interestingly, unlike the thin accretion disk model, in the thick disk model, the gravitational lensing effect of radiation from outside of the equatorial plane causes polarization vectors to spread across the entire image plane, leading to more complex polarization patterns.